# Numerical investigation of buoyancy-aided mixed convective flow past a square cylinder inclined at 45 degrees

### v1

@@ IV Conclusions
In this study, we conducted a detailed examination of buoyancy-aided mixed convective flow past a square cylinder at an angle of incidence $\alpha=45^{\circ}$ for Re $=100$, $0.0\leq Ri\leq 1.0$ and Pr $=0.7$ by performing direct numerical simulations (DNS) using the finite-difference Marker and Cell (MAC) method. We evaluated different integral parameters, such as drag and lift coefficients and Nusselt number, for various Ri and observed that the time-averaged integral parameters increase with Ri, while the RMS integral parameters decrease with Ri. Using Stuart-Landau analysis, we found that the critical Ri for the suppression of vortex shedding in the near-field is 0.681, which matches that predicted by DNS excellently.
At a critical Ri of 0.68, vortex shedding suppression takes place, leading to a steady flow in the near-field, while simultaneously, the far-field develops large-scale unsteadiness similar to a plume, with no range of Ri for which the entire flow field is seen to be stable. At low Ri, the downstream flow behavior resembles that of a wake, having a deficit of momentum in the time-averaged velocity. As Ri is increased, the flow behavior gradually changes into that of a plume, which behaves like a jet, having an excess of momentum in the time-averaged velocity. However, the flow retains characteristics of a wake and shows dual wake-plume behavior in this intermediate Ri range. An inversion of vorticity takes place behind the cylinder for all positive Ri, due to the interplay between the forced and natural convective components of the flow which induce vorticities of opposite nature.
These findings provide valuable insight into the near-field and far-field dynamics of the buoyancy-aided mixed convective flows past cylinders. The suppression of vortex shedding, inversion of vorticity, and onset of the far-field plume-like unsteadiness contribute to our understanding of the characteristics of such flows.
@@ Acknowledgements.
Two of the authors (Kavin Kabilan and Swapnil Sen) would like to thank SURGE, Indian Institute of Technology Kanpur for providing financial support to carry out this work. All authors thank the Indian Institute of Technology Kanpur for providing the computational facilities required to carry out this work.

### v3

@@ IV Conclusions
In this study, we conducted a detailed examination of buoyancy-aided mixed convective flow past a square cylinder at an angle of incidence, $\alpha=45^{\circ}$ for Re $=100$, $0.0\leq Ri\leq 1.0$ and Pr $=0.7$ by performing direct numerical simulations (DNS) using the finite-difference based Marker and Cell (MAC) method.
At a critical Ri between 0.6 and 0.7, vortex shedding suppression takes place, leading to a steady flow in the near-field. Simultaneously, the far-field develops large-scale unsteadiness similar to a plume. There is no range of Ri for which the entire flow field is seen to be steady. At low Ri, the downstream flow behavior resembles that of a wake, having a momentum deficit in the time-averaged velocity. As Ri is increased, the flow behavior gradually transforms into a jet-like plume, having a momentum excess in the time-averaged velocity. The flow retains the characteristics of both a wake and a plume in this intermediate Ri range. An inversion of vorticity takes place behind the cylinder for all positive Ri, because of the interplay between the forced and natural convective components of the flow. These findings provide valuable insight into the near-field and far-field dynamics of the buoyancy-aided mixed convective flows past cylinders.
We evaluated different integral parameters, such as drag and lift coefficients and Nusselt number. The time-averaged parameters ($\overline{C}_{D}$ and $\overline{Nu}$) show an increase with Ri, while the corresponding RMS parameters show a decrease with Ri. Additionally, the inversion length decreases with Ri, while the recirculation length increases with increasing Ri till the critical Ri and decreases subsequently.
@@ Acknowledgements.
Two of the authors (Kavin Kabilan and Swapnil Sen) would like to thank SURGE, Indian Institute of Technology Kanpur, for providing financial support to carry out this work. The authors thank the Indian Institute of Technology Kanpur for providing the computational facilities required to carry out this work.